OpenStudy (anonymous):

once again, i need help please! (cos x)(cot x) + (cos x) = 0 solve over 0, 2pi

OpenStudy (anonymous):

okay, so far i got this: (cos x)(cot x) + cos x = 0 cos x * cot x = - cos x cot x = -cos x/cos x cot x = -1 am I doing this right so far? XD

OpenStudy (kainui):

Lookin' good!

OpenStudy (anonymous):

so, if i did this right, it means that 135 degrees and 315 degrees angles, or 3pi/4 and 7pi/4, is the answer? because tanget is -1, so cot would flip -1/1 so it'd still be -1, right?

hartnn (hartnn):

(cos x)(cot x) + cos x = 0 cos x (cot x +1) = 0 cos x = 0 and cot x+1 =0 you need to consider both of these equations :)

hartnn (hartnn):

because your original equation will get satisfied even when cos x =0, right ?

OpenStudy (anonymous):

cos x = 0, it doesn't work. becuase that means itll be 90 and 270 respectively tan of 90 is undefined, and cot is 0 tan of 270 is undefined and cot is 0 still. thnaks ofr hte help though, it made me second guess myself

hartnn (hartnn):

cos 90 times cot 90 is still 0 so they both do work :)

OpenStudy (anonymous):

my instructor told me i have to idenify the variables by themselves, so i have to evaluate cot x = -1 and cos x = 0, individually.

OpenStudy (anonymous):

\[ cosx (cosx/sinx)+cosx=0 <=> cos^{2}x+cosxsinx=0 <=> cosx(cosx+sinx)=0 \] (i) \[ <=> cosx=0 <=> cosx=cosπ/2 <=> x=2kπ\pmπ/2 \] but \[0\le x\le 2π <=>0\le 2kπ\pmπ/2\le2π <=> 0\le k \pm1/4 \le1\] \[ <=> \mp1/4\le k\le1\mp1/4\] but \[k \in \mathbb{Z}\] so k=0 or k=1 \[x=\pmπ/2 \] or \[x=5π/2\] or \[x=3π/2\] and (ii)\[ cosx +sinx=0\] \[ cosx=-sinx<=> 1=-tanx (*) <=>tanx=-1 <=> tanx=tan7π/4 \] \[<=>x=kπ+7π/4\] but \[x \in [0,2π] \] \[0\le kπ+7π/4\le 2π <=> -7/4\le k\le1/4\] so k=-1,0 k=-1, x=3π/4 k=0, x=7π/4 (*) \[cosx \neq 0 \] so (i) step is cancelled So x=3π/4 and x=7π/4